# Equivariant Spherical Transformer for Efficient Molecular Modeling

### v2

@@ Abstract
SE(3)-equivariant Graph Neural Networks (GNNs) have significantly advanced molecular system modeling by employing group representations. However, their message passing processes, which rely on tensor product-based convolutions, are limited by insufficient non-linearity and incomplete group representations, thereby restricting expressiveness. To overcome these limitations, we introduce the Equivariant Spherical Transformer (EST), a novel framework that leverages a Transformer structure within the spatial domain of group representations after Fourier transform. We theoretically and empirically demonstrate that EST can encompass the function space of tensor products while achieving superior expressiveness. Furthermore, EST’s equivariant inductive bias is guaranteed through a uniform sampling strategy for the Fourier transform. Our experiments demonstrate state-of-the-art performance by EST on various molecular benchmarks, including OC20 and QM9.
@@ 1 Introduction
Graph neural networks (GNNs) have become increasingly prevalent for modeling molecular systems and approximating quantum mechanical calculations, providing crucial support for various computational chemistry tasks, including drug discovery senior2020improved and material design zitnick2020introduction . Compared to traditional calculation methods like Density Functional Theory (DFT), GNNs significantly reduce the computational cost of quantum property prediction from hours or days to fractions of a second. However, directly applying regular GNNs (e.g. GCN kipf2016semi and GIN xu2018powerful ) to 3D molecular conformations often yields poor results due to their neglect of inherent physical constraints. To overcome these limitations, equivariant GNNs have emerged as a more promising avenue, effectively capturing intricate atomic interactions and addressing multiple crucial challenges.
@@
In molecular systems, SE(3)-invariance and SE(3)-equivariance are fundamental constraints. For instance, molecular energies remain constant under rotation, while atomic forces rotate in concert with the molecule equivariantly. Early invariant models, such as SchNet schutt2018schnet and HIP-NN lubbers2018hierarchical , rely on interatomic distances in their message-passing blocks, consequently limiting their ability to capture interactions involving triplets or quadruplets of atoms miller2020relevance . Directional GNNs gasteiger2020directional ; gasteiger2021gemnet address this problem by explicitly incorporating bond angles and dihedral angles. Despite performance improvements, their expressiveness remained constrained as they primarily extracted features from invariant and handcrafted features. To capture deeper features, SE(3)-equivariant GNNs employ group representations as node embeddings and construct steerable message-passing blocks thomas2018tensor ; brandstetter2021geometric ; liao2023equiformer , where tensor product are applied to capture equivariant interactions between group embeddings. These GNNs achieve improved performance using only raw molecular structures, without providing handcrafted features. Nevertheless, tensor product operations inherently suffer from limited non-linearity thomas2018tensor , and their expressiveness is bounded by the degree of group representations dym2021on ; joshi2023expressive . An alternative equivariant approach processes features in the spatial domain after Fourier transforming from group representations, where embeddings are defined by square-integrable spherical functions (see Figure 1(a)). Prior works zitnick2022spherical ; passaro2023reducing ; liao2024equiformerv typically apply simple point-wise neural networks to model signals on these spherical functions, operating independently on each orientation (see Figure 1(b)). Introducing nonlinearity across different orientations within these spherical functions is rarely considered due to the risk of violating equivariance. Nevertheless, we show that this operation can be performed equivariantly, potentially providing additional nonlinearity and expressiveness (Further analysis is provided in Section 3.2.)
In this paper, we present the Equivariant Spherical Transformer (EST), a novel message-passing framework for accurate atomic interaction modeling. First, we use the spherical functions in the spatial domain as node and edge embeddings, and transform it to point sequences, where each point represent a unique orientation on the spherical functions. A spherical attention mechanism is then introduced to capture dependencies within these point sequences (see Figure 1(c)). Furthermore, EST incorporates a mixture of hybrid experts structure, utilizing feed-forward networks in both spatial domain and original harmonic domain to effectively balance equivariance and model capacity. The equivariant inductive bias of EST is guaranteed through our implementation of a uniform sampling strategy in spatial domain. Our experiments show that EST-based architectures achieve state-of-the-art (SOTA) performance on various benchmarks such as OC20 chanussot2021open and QM9 ramakrishnan2014quantum , exhibiting particularly stable and significant improvements for complex molecular systems.
@@
Our contributions are: (i) We propose EST, a novel framework whose expressiveness provably subsumes traditional tensor products. (ii) By employing an uniform sampling implementation, we guarantee the equivariance of EST. (iii) We conduct a range of experiments for quantum property prediction tasks, demonstrating superior performance of EST. Moreover, our ablation studies confirming the enhanced expressiveness and equivariance of EST.
2 Related Work and Preliminaries
@@
In this section, we discuss related work and the corresponding mathematical background relevant to equivariant GNNs. We begin by listing the notations frequently used throughout the paper. We denote the unit sphere as $\mathbb{S}^{2}$, where the coordinate of a spherical point (or orientation) $\vec{\mathbf{p}}=(\theta,\varphi)$ is represented by its polar angle $\theta$ and azimuth angle $\varphi$. The symbol $\mathbb{R}$ represents the set of real numbers, while $\mathbf{R}$ denotes a rotation matrix for 3D vectors. We use $\cdot$$\cdot$] to indicate tensor concatenation and $\tilde{\cdot}$ to represent the update of tensors, respectively.
Figure 1: Overview of operations in the spherical spatial domain. (a) Applying the Fourier transform to group steerable representations, projecting them onto the spatial domain and storing them via spherical sampling. (b) A conventional point-wise operation on the sphere, where a function $g(\cdot)$ combines features across different spherical functions at a given orientation. (c) The proposed EST operation, where $g(\cdot)$ represents a Transformer framework designed to model global dependencies and interactions among different orientations on the sphere.
@@
Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with nodes $v_{i}\in\mathcal{V}$ and edges $e_{ij}\in\mathcal{E}$. Each node $v_{i}$ has an embedding $\mathbf{x}_{i}$ and an attribute $\mathbf{z}_{i}$, and each edge $e_{ij}$ has an attribute $\mathbf{a}_{ij}$. The Message Passing Neural Network (MPNN) schutt2018schnet , a specific type of GNN, updates node embeddings through a message block $\mathbf{M}(\cdot)$ and an update block $\mathbf{U}(\cdot)$ via the following steps:
@@
Spherical harmonics, a set of orthonormal basis functions defined over the sphere $\mathbb{S}^{2}$, are commonly employed in equivariant models. The real-valued spherical harmonics are typically denoted as $\{Y^{(l,m)}:\mathbb{S}^{2}\to\mathbb{R}\}$, where $l$ represents the degree and $m$ the order. For any orientation $\vec{\mathbf{p}}$, we define $\mathbf{Y}^{(l)}(\vec{\mathbf{p}})=[Y^{(l,-l)}(\vec{\mathbf{p}}),Y^{(l,-l+1)}(\vec{\mathbf{p}}),...,Y^{(l,l)}(\vec{\mathbf{p}})]$, a vector of size $2l+1$, and $\mathbf{Y}^{(0\to l)}(\vec{\mathbf{p}})=[\mathbf{Y}^{(0)}(\vec{\mathbf{p}}),...,\mathbf{Y}^{(l)}(\vec{\mathbf{p}})]$, a vector of size $(l+1)^{2}$.
@@
$\mathbf{Y}^{(l)}(\mathbf{R}\vec{\mathbf{p}})=\mathbf{D}^{(l)}(\mathbf{R})\mathbf{Y}^{(l)}(\vec{\mathbf{p}}),$ (3)
where $\mathbf{D}^{(l)}(\mathbf{R})$ is a $(2l+1)\times(2l+1)$ matrix known as the Wigner-D matrix of degree $l$. Notably, $\mathbf{D}^{(1)}(\mathbf{R})=\mathbf{R}$. Thus, $\mathbf{R}$ and $\mathbf{D}^{(l)}(\mathbf{R})$ can represent $T(g)$ and $T^{\prime}(g)$ in Equation (2). Following the convention in (chami2019hyperbolic, ; brandstetter2021geometric, ), we say that $\mathbf{Y}^{(l)}(\vec{\mathbf{p}})$ is steerable by the Wigner-D matrix of the same degree $l$. Furthermore, a vector that transforms according to an $l$-degree Wigner-D matrix is termed an $l$-degree steerable vector or a type-$l$ vector, residing in the vector space $\mathbb{V}_{l}$. Further mathematical details are available in Appendix A.1.
A common practice in equivariant models is to use steerable vectors as node embedding and encode relative positions $\vec{\mathbf{r}}_{ij}$ using spherical harmonics as edge attributes. TFN thomas2018tensor and NequIP batzner2022 provide a general framework that learns the interaction between node embeddings and edge attributes through equivariant convolution filters. These filters are composed of equivariant operations such as degree-wise linear (DW-Linear) layers, Clebsch-Gordan (CG) tensor products, and Gate mechanisms. SEGNN (brandstetter2021geometric, ) extends this convolution framework by more non-linear operations to achieve enhanced learning ability. SE(3)-Transformer fuchs2020se and Equiformer liao2023equiformer introduce an attention mechanism that assigns rotation-invariant weights to edge attributes, thereby improving the learning of key atomic interactions. Nevertheless, the core operation in these works remains the convolution filter based on the CG tensor product. Additionally, GotenNet aykentgotennet proposes an effective and lightweight structure that separates steerable vectors into an $l=0$ invariant component and $l>0$ equivariant components, subsequently interacting these parts using inner products and scalar multiplication. In summary, due to the requirement of equivariance, most equivariant model architectures are constrained to specific equivariant operations. In this work, we transform the steerable vector into spherical spatial domain, enabling the application of more flexible and highly non-linear structures within message passing.
@@
where $d\Omega$ denotes the unit sphere and $Y^{(l,m)*}(\cdot)$ is the conjugate of the spherical harmonic function. The function $f(\vec{\mathbf{p}})$ and its coefficients $\mathbf{x}$ are referred to as the representations in the spherical spatial domain and the harmonic domain, respectively. SCN zitnick2022spherical introduced an approach that applies a simple point-wise neural network in the spherical spatial domain. This technique has been subsequently adopted by several works passaro2023reducing ; liao2024equiformerv within their message-passing frameworks. However, point-wise neural networks inherently limit their expressiveness. Our work aims to extend the capabilities of learning within the whole spherical spatial domain.
@@ 3.1 Review of Clebsch-Gordan Tensor Product
where $\mathcal{C}^{(l,m)}_{(l_{1},m_{1})(l_{2},m_{2})}$ are the CG coefficients, a sparse tensor yielding non-zero terms when $|l_{1}-l_{2}|\leq l\leq(l_{1}+l_{2})$. These products enable interactions between various combinations of $l_{1}$ and $l_{2}$, crucial for the expressive power of MPNNs in representing latent equivariant functions dym2021on ; joshi2023expressive . However, CG tensor products lack high non-linearities brandstetter2021geometric , often requiring the stacking of multiple layers to capture complex molecular features. Furthermore, their high computational complexity ($\mathcal{O}(L^{6})$) luo2024enabling makes architectures with stacked tensor product layers computationally expensive brandstetter2021geometric . While techniques like Gate mechanisms brandstetter2021geometric ; liao2023equiformer and geometry-aware tensor attention aykent2025gotennet can be used to mitigate this limitation, they apply non-linear transformations to the invariant ($l=0$) representations. Higher-degree representations ($l>0$) are then updated through a scaling multiplication:
@@
where $C_{inv}$ is an invariant scalar, $\mathrm{NL}(\cdot)$ is a network with non-linear functions, and $\mathrm{d}(\cdot):\mathbb{V}_{l}\rightarrow\mathbb{V}_{0}$ maps equivariant features to invariants, such as an inner product. This indirect exchange of information between higher-degree representations still limits the overall expressiveness.
@@ 3.2.1 Node embedding
For representation with steerable vectors, we first define a maximal degree $L$. Each node $n$’s embedding $\mathbf{x}_{n}$ comprises $C$ channels, and each channel is a concatenation of steerable vectors from degree 0 to $L$: $\mathbf{x}_{n,c}=[x^{(0)}_{n,c},\mathbf{x}^{(1)}_{n,c},...,\mathbf{x}^{(L)}_{n,c}$]. As a result, the dimension of a steerable node embedding is $(L+1)^{2}\times C$.
Spatial Representation.
Steerable representations $\mathbf{x}_{n,c}$ can be transformed into a spherical function $f_{n,c}(\vec{\mathbf{p}})$ via Equation (4), with the summation truncated at a maximum degree $L$. Furthermore, $f_{n,c}(\vec{\mathbf{p}})$ is discretely represented by sampling $S$ points on the sphere, yielding a spatial node embedding $\mathbf{f}^{*}_{n}\in\mathbb{R}^{S\times C}$. It can be expressed as the concatenation over sampled points:
@@
We use the term $\mathbf{f}^{*}_{n,c,s}$ to denote the signal at the sampled point $\vec{\mathbf{p}}_{s}$ of channel $c$ of node $n$. The inverse transform in Equation (5) can convert $\mathbf{f}^{*}_{n}$ back to $\mathbf{x}_{n}$.
In our model, the standard state of the node embedding is a steerable representation and it is transformed into a spatial representation when specific spherical operations are performed. Furthermore, the initial node embedding is constructed from two components: invariant mappings of the atomic attributes (the $l=0$ part) and the spherical harmonics of its neighbors’ relative positions (the $l>0$ part).
Operations in spatial domain
[…]

### v3

@@ Abstract
Equivariant Graph Neural Networks (GNNs) have significantly advanced the modeling of 3D molecular structure by leveraging group representations. However, their message passing, heavily relying on Clebsch-Gordan tensor product convolutions, suffers from restricted expressiveness due to the limited non-linearity and low degree of group representations. To overcome this, we introduce the Equivariant Spherical Transformer (EST), a novel plug-and-play framework that applies a Transformer-like architecture to the Fourier spatial domain of group representations. EST achieves higher expressiveness than conventional models while preserving the crucial equivariant inductive bias through a uniform sampling strategy of spherical Fourier transforms. As demonstrated by our experiments on challenging benchmarks like OC20 and QM9, EST-based models achieve state-of-the-art performance. For the complex molecular systems within OC20, small models empowered by EST can outperform some larger models and those using additional data. In addition to demonstrating such strong expressiveness,we provide both theoretical and experimental validation of EST’s equivariance as well, paving the way for new research in this area.
@@ 1 Introduction
Graph neural networks (GNNs) are increasingly used for modeling molecular systems and approximating quantum mechanical calculations, providing crucial support for computational chemistry tasks like drug discovery [1] and material design [2]. Compared to traditional methods such as Density Functional Theory (DFT), GNNs can predict quantum properties in fractions of a second, significantly reducing computational cost from hours or days. Among the various GNN architectures, SE(3)-equivariant GNN [3, 4] has become a leading variant due to their inherent physical constraints, which can effectively capture intricate atomic interactions and address key challenges in molecular modeling.
Figure 1: Overview of operations in the spherical spatial domain. (a) Examples of group representation. A Fourier transform is applied to project steerable vectors onto the spatial domain, which are then stored through sampling (i.e., discretization). (b) Tensor product operation. Steerable vectors are combined using the Clebsch-Gordan coefficient $\mathcal{C}$. (c) Conventional point-wise operation. All channels within the same orientation are processed by a vanilla neural network ((with linear projection $W$ and activation $\sigma(\cdot)$), and no interaction is performed across different orientations. (d) Proposed EST operation. A Transformer-based structure captures dependencies across all orientation pairs to model complex atomic behavior.
@@
In molecular systems, SE(3)-invariance and SE(3)-equivariance are fundamental constraints. For instance, molecular energies remain constant under rotation, while atomic forces rotate in concert with the molecule equivariantly. Early invariant models, such as SchNet [5] and HIP-NN [6], rely on interatomic distances in their message-passing blocks, consequently limiting their ability to capture interactions involving triplets or quadruplets of atoms [7]. Directional GNNs [8, 9] address this problem by explicitly incorporating bond angles and dihedral angles. While these models showed performance improvements, their expressiveness remains constrained due to a heavy reliance on handcrafted features. To capture deeper features, SE(3)-equivariant GNNs employ group representations as node embeddings and construct steerable message-passing blocks [3, 4, 10], where tensor product are applied to capture equivariant interactions between group embeddings (see Figure 1(b)). These GNNs achieve improved performance through end-to-end modeling, eliminating the need for hand-crafted features. Nevertheless, tensor product operations inherently suffer from limited non-linearity [3], and their expressiveness is bounded by the degree of group representations [11, 12, 13]. An alternative equivariant approach is to capture features in the spatial domain by applying a Fourier Transform (FT) to group representations, where embeddings are defined by square-integrable spherical functions (see the right side of Figure 1(a)). Prior works [14, 15, 16] typically apply simple point-wise neural networks to extract features from these spherical functions, operating independently on each orientation (see Figure 1(c)). However, operations introducing non-linearity across different orientations in the spatial domain are rarely considered, as they risk violating equivariance. We demonstrate that this challenge can be overcome by subtly redesigning the FT sampling and architecture, which can also provide additional expressiveness (Further analysis is provided in Section 3.2).
This paper presents the Equivariant Spherical Transformer (EST), a plug-and-play equivariant framework for high-fidelity modeling of atomic interactions. Our methodology initiates the process in the Fourier spatial domain of group representations by converting the spatial representation into a sequence of points. Subsequently, a Transformer-based architecture, as depicted in Figure 1(d), is employed to capture dependencies across different orientations on this sequence. A key finding of our work is that conventional spherical FT sampling methods, exemplified by the e3nn grid [17], significantly degrade equivariance. To mitigate this issue, we introduce a new, provably spherical sampling strategy grounded in the Fibonacci lattice, which we validate through theoretical and empirical analyses. Additionally, we incorporate a “hybrid mixture of experts” structure within our Transformer, featuring feed-forward networks in both the spatial and harmonic domains to strike an optimal balance between equivariance and model capacity. Experiments demonstrate that EST-based architectures attain state-of-the-art (SOTA) performance on standard benchmarks such as OC20 [18] and QM9 [19], with stable and notable improvements. On the complex molecular systems of the OC20 dataset, a compact EST model surpasses the performance of larger models and those leveraging external training data, thereby attesting to its strong expressiveness for modeling molecular systems.
@@
Our contributions are: (i) We propose EST, a plug-and-play framework whose expressiveness subsumes that of traditional tensor products-based operations. (ii) We guarantee the equivariance of EST by subtly employing an approximately uniform sampling implementation on spherical FT. (iii) We demonstrate the superior performance of EST through extensive experiments. Furthermore, our ablation studies confirm the equivariance and enhanced expressiveness of EST, establishing it as a promising building block for a wide range of equivariant architectures.
2 Preliminaries and Related Work
@@
In this section, we discuss related work and the corresponding mathematical background relevant to equivariant GNNs. We begin by listing the notations frequently used throughout the paper. We denote the unit sphere as $\mathbb{S}^{2}$, where the coordinate of a spherical point (or orientation) $\vec{\mathbf{p}}=(\theta,\varphi)$ is represented by its polar angle $\theta$ and azimuth angle $\varphi$. The symbol $\mathbb{R}$ represents the set of real numbers, while $\mathbf{R}$ denotes a rotation matrix for 3D vectors. We use $\cdot$$\cdot$] or $\oplus$ to indicate tensor concatenation and $\tilde{\cdot}$ to represent the update of tensors, respectively.
@@
Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with nodes $v_{i}\in\mathcal{V}$ and edges $e_{ij}\in\mathcal{E}$. Each node $v_{i}$ has an embedding $\mathbf{x}_{i}$ and an attribute $\mathbf{z}_{i}$, and each edge $e_{ij}$ has an attribute $\mathbf{a}_{ij}$. The Message Passing Neural Network (MPNN) [5], a specific type of GNN, updates node embeddings through a message block $\mathbf{M}(\cdot)$ and an update block $\mathbf{U}(\cdot)$ via the following steps:
@@
Spherical harmonics, a set of orthonormal basis functions defined over the sphere $\mathbb{S}^{2}$, are commonly employed in equivariant models. The real-valued spherical harmonics are typically denoted as $\{Y^{(l,m)}:\mathbb{S}^{2}\to\mathbb{R}\}$, where $l$ represents the degree and $m$ represents the order. For any orientation $\vec{\mathbf{p}}$, we define $\mathbf{Y}^{(l)}(\vec{\mathbf{p}})=[Y^{(l,-l)}(\vec{\mathbf{p}}),Y^{(l,-l+1)}(\vec{\mathbf{p}}),...,Y^{(l,l)}(\vec{\mathbf{p}})]$, a vector of size $2l+1$, and $\mathbf{Y}^{(0\to l)}(\vec{\mathbf{p}})=[\mathbf{Y}^{(0)}(\vec{\mathbf{p}}),...,\mathbf{Y}^{(l)}(\vec{\mathbf{p}})]$, a vector of size $(l+1)^{2}$.
@@
$\mathbf{Y}^{(l)}(\mathbf{R}\vec{\mathbf{r}})=\mathbf{D}^{(l)}(\mathbf{R})\mathbf{Y}^{(l)}(\vec{\mathbf{r}}),$ (3)
where $\mathbf{D}^{(l)}(\mathbf{R})$ is a $(2l+1)\times(2l+1)$ matrix known as the Wigner-D matrix of degree $l$. Notably, $\mathbf{D}^{(1)}(\mathbf{R})=\mathbf{R}$. Thus, $\mathbf{R}$ and $\mathbf{D}^{(l)}(\mathbf{R})$ can represent $T(g)$ and $T^{\prime}(g)$ in (2). Following the convention in [20, 4], we say that $\mathbf{Y}^{(l)}(\vec{\mathbf{p}})$ is steerable by the Wigner-D matrix of the same degree $l$. Furthermore, a vector that transforms according to an $l$-degree Wigner-D matrix is termed an $l$-degree steerable vector or a type-$l$ vector, residing in the vector space $\mathbb{V}_{l}$. Further mathematical details are available in Appendix A.1.
Equivariant architectures
A common practice in equivariant models is to use steerable vectors as node embedding and encode relative positions $\vec{\mathbf{r}}_{ij}$ using spherical harmonics as edge attributes. TFN [3] and NequIP [21] provide a general framework that learns the interaction between node embeddings and edge attributes through equivariant convolution filters. These filters are composed of equivariant operations such as degree-wise linear (DW-Linear) layers, Clebsch-Gordan (CG) tensor products, and Gate mechanisms. SEGNN [4] extends this convolution framework by more non-linear operations to achieve enhanced learning ability. SE(3)-Transformer [22] and Equiformer [10] introduce an attention mechanism that assigns rotation-invariant weights to edge attributes, thereby improving the learning of key atomic interactions. Nevertheless, the core operation in these works remains the convolution filter based on the CG tensor product. Additionally, GotenNet [23] proposes an effective and lightweight structure that separates steerable vectors into an $l=0$ invariant component and $l>0$ equivariant components, subsequently interacting these parts using inner products and scalar multiplication. In summary, due to the requirement of equivariance, most equivariant model architectures are constrained to specific equivariant operations. In this work, we transform the steerable vector into spherical spatial domain, enabling the application of more flexible and highly non-linear structures within message passing.
@@
where $d\Omega$ denotes the unit sphere and $Y^{(l,m)*}(\cdot)$ is the conjugate of the spherical harmonic function. The function $f(\vec{\mathbf{p}})$ and its coefficients $\mathbf{x}$ are referred to as the representations in the spherical spatial domain and the harmonic domain, respectively. These processes are illustrated in Figure 1a. SCN [14] introduced an approach that applies a simple point-wise neural network in the spherical spatial domain. This technique has been subsequently adopted by several works [15, 16] within their message-passing frameworks. However, point-wise neural networks inherently limit their expressiveness. Our work aims to extend the capabilities of learning within the whole spherical spatial domain.
@@ 3.1 Review of Clebsch-Gordan Tensor Product
where $\mathcal{C}^{(l,m)}_{(l_{1},m_{1})(l_{2},m_{2})}$ are the CG coefficients, a sparse tensor yielding non-zero terms when $|l_{1}-l_{2}|\leq l\leq(l_{1}+l_{2})$. These products enable interactions between various combinations of $l_{1}$ and $l_{2}$, crucial for the expressive power of MPNNs in representing latent equivariant functions [11, 12]. However, CG tensor products lack high non-linearities [4], often requiring the stacking of multiple layers to capture complex molecular features. Furthermore, their high computational complexity ($\mathcal{O}(L^{6})$) [24] makes architectures with stacked tensor product layers computationally expensive [4]. While techniques like Gate mechanisms [4, 10] and geometry-aware tensor attention [25] can be used to mitigate this limitation, they apply non-linear transformations to the invariant ($l=0$) representations. Higher-degree representations ($l>0$) are then updated through a unified scaling multiplication:
@@
where $C_{inv}$ is an invariant scalar, $\mathrm{NL}(\cdot)$ is a non-linear network, and $\mathrm{d}(\cdot):\mathbb{V}_{l}\rightarrow\mathbb{V}_{0}$ maps equivariant features to invariants, such as an inner product. This indirect exchange of information between higher-degree ($l>0$) representations still limits the overall expressiveness.
@@ 3.2.1 Node embedding
For representation with steerable vectors, we first define a maximal degree $L$. Each node $n$’s embedding $\mathbf{x}_{n}$ comprises $C$ channels, and each channel $c$ is a concatenation of steerable vectors from degree 0 to $L$: $\mathbf{x}^{(0\to L)}_{n,c}=[x^{(0)}_{n,c},\mathbf{x}^{(1)}_{n,c},...,\mathbf{x}^{(L)}_{n,c}$]. As a result, the dimension of a steerable node embedding is $(L+1)^{2}\times C$.
Spatial Representation
Steerable representations $\mathbf{x}_{n,c}$ can be transformed into a spherical function $f_{n,c}(\vec{\mathbf{p}})$ via (4), with the summation truncated at a maximum degree $L$. Furthermore, $f_{n,c}(\vec{\mathbf{p}})$ is discretely represented by sampling $S$ points on the sphere, yielding a spatial node embedding $\mathbf{f}^{*}_{n}\in\mathbb{R}^{S\times C}$. It can be expressed as the concatenation over sampled points:
@@
We use the term $\mathbf{f}^{*}_{n,c,s}$ to denote the signal at the sampled point $\vec{\mathbf{p}}_{s}$ of channel $c$ of node $n$. The inverse transform in (5) can convert $\mathbf{f}^{*}_{n}$ back to $\mathbf{x}_{n}$.
The steerable and spatial representations can be converted into each other. This property allows our proposed EST to be easily integrated as a plug-and-play module into a variety of equivariant models.
The significance of operating in the spatial domain
In the spatial representation, the explicit degree and order are no longer directly accessible. Instead, higher-degree information is encoded as geometric features distributed across the spherical surface, as illustrated in Figure 1(a). Intuitively, by modeling the interdependencies of these geometric features, the model can effectively exchange information between implicit degrees, thereby approximating the capability of the tensor product discussed in Section 3.1. However, prior methods [14, 15, 16] only consider the feature dependencies across channels within the same sampling point (e.g. $\mathbf{f}^{*}_{c_{1},s_{1}},\mathbf{f}^{*}_{c_{2},s_{1}},...,\mathbf{f}^{*}_{c_{C},s_{1}}]$), as depicted in Figure 1(c), overlooking geometric features composed of multiple points. In contrast, our proposed EST can model dependencies across various sampling points (e.g. $\mathbf{f}^{*}_{s_{1}},\mathbf{f}^{*}_{s_{2}},...\mathbf{f}^{*}_{s_{S}}]$). We make two key claims: (i) our EST is SO(3)-equivariant by an uniform FL sampling strategy on the sphere, and (ii) its expressive power can surpass that of tensor product-based frameworks.